OpenStudy (sharkfinsoup):

h(x)=(g o f)(x)=1/(x+3)^2 which of the following could be a possible decomposition of h(x)? A. f(x)=1/x^2;g(x)=x+3 B. f(x)=x+3; g(x)=1/x^2 C.f(x)=x^2; g(x)=x+3 D. f(x)=1/x;g(x)=x+3

OpenStudy (sharkfinsoup):

I think its C f(x)=x+3; g(x)=1/x^2 Because if you take f(x)=x+3 g(x)=1/x^2 plug the f(x) into the g(x) formula this is what you get g(x)=1/x^2 g(x)=1/(x+3)^2. I think

OpenStudy (unklerhaukus):

Do you mean B or C?

OpenStudy (sharkfinsoup):

oh sorry I meant B

OpenStudy (unklerhaukus):

yeah, f(x) = x+3, g(x) = 1/x^2 (g o f) (x) = 1/(x+3)^2.

OpenStudy (unklerhaukus):

What would you get with option A?

OpenStudy (sharkfinsoup):

wouldn't that be (f o g)(x)

OpenStudy (unklerhaukus):

with A: f(x)=1/x^2; g(x)=x+3 fog would work to get the right result: (f o g)(x) = f(x+3) = 1/(x+3)^2 But what would gof give?

OpenStudy (unklerhaukus):

(gof)(x) = g( f(x) ) = g( 1/x^2 ) = ?

OpenStudy (sharkfinsoup):

wait so it would be (1/x^2)+3

OpenStudy (unklerhaukus):

that's correct

OpenStudy (unklerhaukus):

As we can now see B is the only correct option.

OpenStudy (sharkfinsoup):

Right that's great. Thank you so much :)

OpenStudy (unklerhaukus):

\[\ddot\smile\]